A Carnot refrigerator is operated between two heat reservoirs at temperatures of 320 K

and 270 K.

  • If in each cycle the refrigerator receives 415 J of heat energy from the reservoir at 270 K, how many joules of heat energy does it deliver to the reservoir at 320 K?
  • If the refrigerator completes 165 cycles each minute, what power input is required to operate it?
  • What is the coefficient of performance of the refrigerator?
1
Expert's answer
2021-02-05T17:06:31-0500

The work input to a refrigerator is given by the relation

"W = |Q_H|-|Q_C|" (1)

For a Carnot refrigerator |QH| and |QC| are related to T1 and T2 by the formula

"\frac{|Q_H|}{|Q_C|} = \frac{T_H}{T_C}" (2)

The coefficient of performance of a Carnot refrigerator is

"k = \frac{T_C}{T_H-T_C}" (3)

TH = 320 K

TC = 270 K

|QC| = 415 J

The amount of heat delivered to the high temperature reservoir is, by using equation (2), given by

"|Q_H| = |Q_C| \frac{T_H}{T_C} \\

= \frac{415 \times 320}{270} \\

= 492 \;J"

The number of cycles per minute is n = 165 cycles/min

"= \frac{165}{60} \;cycles/s"

Using equation (1), the work input required to run the refrigerator is

"W = |Q_H| -|Q_C| \\

= 492 -415 \\

= 77 \;J"

The power input required to operate the refrigerator for n cycle per second is

"W \times n \;J/s"

Hence, the total input power required

"= \frac{77 \times 165}{60} \;J/s \\

= 212 \;W"

Using equation (3), the coefficient of performance of the Carnot refrigerator is

"k = \frac{270}{320-270} \\

= \frac{270}{50} \\

= 5.4"